# Complex Magnetic Behavior in RuO2 Thin Films: Strain, Surface Effects, and Altermagnetic Signatures

### v1

@@ III.2 RuO2(110) thin films on TiO2(110)
For this system, each layer contains two inequivalent Ru atoms (Ru1 and Ru2), yielding a total of six magnetic sites and ten distinct AFM configurations. We systematically examined all these AFM arrangements. Each configuration was initially imposed using magnetic constraints to stabilize the desired spin pattern; the constraints were subsequently released to allow full self-consistent relaxation. Among all tested configurations, the one initialized with antiparallel magnetic moments between Ru1 and Ru2 within each layer consistently relaxes to the lowest-energy state, characterized by opposite signs of the magnetic moments on Ru1 and Ru2, with incomplete compensation. The remaining AFM initializations either converge to higher-energy solutions or evolve into this same configuration upon relaxation. Moreover, for several slab thicknesses, calculations starting from a ferromagnetic state also converge to this energetically favorable AFM ordering. These results indicate that the ground state generally favors antiparallel alignment between Ru1 and Ru2 within each layer; accordingly, this antiparallel arrangement is adopted as the initial magnetic configuration in all subsequent calculations.
@@
The only deviation from this trend occurs in the 6-L slab, where the Ru atoms in the third layer develop magnetic moments of the same sign on the two sublattices, despite being initialized in the constrained AFM configuration. Nevertheless, the magnetization in this layer remains strongly uncompensated, since the moment on Ru2 is significantly smaller than that on Ru1, resulting in a ferromagnetic alignment of the two sublattices, albeit with strongly unequal magnetic moments. The magnetic moments in the third layer are 0.33 $\mu\mathrm{B}$ on Ru1 and 0.04 $\mu\mathrm{B}$ on Ru2 for the bridge configuration, whereas for the cus configuration the corresponding values are 0.30 $\mu\mathrm{B}$ and 0.007 $\mu\mathrm{B}$, respectively.
@@
Although full structural relaxations were carried out for all slab geometries, Fig. 4 shows that atomic relaxation has only a minor effect on the magnetic moments of the deeper RuO2(110) layers. Moreover, when the magnetic moment of the surface layer is artificially enhanced using a Hubbard $U$ correction, the resulting increase remains highly localized: the magnetic moments in the subsurface and deeper layers change only negligibly. This behavior demonstrates that the magnetization generated at the surface due to symmetry breaking [20] does not propagate into the deeper layers. Instead, the magnetic moments of the Ru atoms are primarily determined by strain-induced changes in their local atomic environment arising from interfacial tensile strain, rather than by long-range induction from the surface.
@@
Figure 3: Magnetic moments of Ru1 and Ru2 as a function of layer index and film thickness. Filled and open symbols denote Ru1 and Ru2, respectively, for RuO2(110) thin films on a TiO2(110) substrate, with thicknesses ranging from 1-L to 7-L. For each thickness, magnetic moments are shown for all Ru layers within the slab. Data corresponding to the same film thickness are connected by lines of the same color to guide the eye.
@@
Figure 4: Effect of atomic relaxation and induced magnetization on Ru magnetic moments in a 5-L RuO2(110) slab on TiO2(110). Magnetic moments of Ru1 (filled symbols) and Ru2 (open symbols) are shown as a function of the Ru layer index.
It is also instructive to compare the electronic structure of a compensated AFM RuO2(110)/TiO2(110) thin film with that of its true magnetic ground state. To this end, we calculate and compare the band structures of a 5-L RuO2(110) slab on a TiO2(110) substrate for two cases: (i) a constrained, fully compensated AFM configuration, which hosts nonrelativistic altermagnetic spin-split bands, and (ii) the unconstrained AFM ground state. In both cases, the non-relaxed geometry is employed to avoid additional effects arising from structural relaxation.
@@
To make the comparison between the two configurations meaningful, we adjusted the fixed magnetic moments of the Ru atoms in the constrained AFM case so that their total absolute magnetization matches that of the unconstrained AFM ground state, where the magnetic moments vary from layer to layer (Fig. 3). Here, the total absolute magnetization is defined as $M_{\mathrm{abs}}=\int_{\mathrm{cell}}\left|n_{\mathrm{up}}(\mathbf{r})-n_{\mathrm{down}}(\mathbf{r})\right|\,d^{3}r$, where $n_{\mathrm{up}}(\mathbf{r})$ and $n_{\mathrm{down}}(\mathbf{r})$ are the spin-up and spin-down charge densities. By testing different values, we found that fixing the Ru moments to approximately $0.3~\mu_{\mathrm{B}}$ reproduces the total absolute magnetization of the unconstrained ground state. This value is therefore used in the constrained AFM calculation.
@@
As shown in Fig. 5, the band structure of the magnetic ground state exhibits a substantially larger nonrelativistic spin splitting than that of the compensated AFM configuration. This enhancement arises because, in the true magnetic ground state, the Ru magnetic moments are not fully compensated, in contrast to the compensated AFM limit, where the net magnetization vanishes.
@@
Figure 5: Nonrelativistic band structures of a 5-L RuO2(110) slab on a TiO2(110) substrate. The left panel shows the band structure of the unconstrained magnetic ground state, while the right panel shows the band structure obtained with Ru magnetic moments constrained to enforce a fully compensated AFM configuration. In both cases, the non-relaxed atomic structure is used. Red and blue lines denote spin-up and spin-down bands, respectively.
III.3 RuO2(100) and (001) thin films on TiO2(100) and (001)
@@
We next consider the remaining low-index RuO2 surfaces, namely the (100) and (001) orientations. Owing to computational constraints, only one film thickness is examined for each case. For both orientations, we study a 12-layer RuO2 film grown on TiO2. Figure 6 shows the layer-resolved magnetic moments of the two inequivalent Ru sites (Ru1 and Ru2), before and after structural relaxation.
@@
For both surfaces, atomic relaxation has a markedly stronger effect than in the (110) case. In the relaxed (001) and (100) films, magnetic moments in the outermost three layers are strongly suppressed, while deeper layers near the RuO2/TiO2 interface develop substantially larger local moments. By contrast, the unrelaxed films exhibit an almost ideal AFM arrangement, with Ru1 and Ru2 moments of similar magnitude and opposite sign; this ordering is significantly weakened upon relaxation.
@@
Despite these differences, all systems, relaxed and unrelaxed, (001), (100), and (110), exhibit a characteristic layer-by-layer oscillation in the magnitude of the local magnetic moments. This behavior can be attributed to the metallic nature of the films: electronic states are confined between the surface–vacuum boundary and the buried interface, forming quasi-two-dimensional states analogous to a particle-in-a-box with finite potential barriers. The resulting quantum interference gives rise to oscillatory magnetic responses across the film thickness.
@@ III.4 Intrinsic magnetism of freestanding RuO2 thin films
To further explore the role of the substrate, we also study freestanding RuO2 thin films with different surface orientations; see Fig. 1. Two classes of films are considered: (i) films constrained to the experimental lattice parameters of rutile TiO2, and (ii) films constrained to the experimental bulk lattice parameters of RuO2. In both cases, only the internal atomic coordinates are relaxed, while the lattice parameters are fixed. Since the in-plane rutile lattice constant of RuO2 is approximately 0.1 Å larger than that of TiO2, constraining the film to the TiO2 lattice imposes in-plane tensile strain. Therefore, adopting the TiO2 lattice parameters provides an effective way to apply biaxial tensile strain to RuO2 thin films.
Figure 7 shows the layer-resolved Ru magnetic moments for the (110), (100), and (001) orientations, both with and without tensile strain. In the unstrained (110) and (100) films, finite magnetic moments are largely confined to the first two surface layers and decay rapidly toward the bulk. The unstrained (110) result is consistent with recent reports (Ref. [20]).
In contrast, the (001) surface converges to a nonmagnetic solution, with no Ru atom exhibiting a discernible magnetic moment. This finding differs from recent reports of finite magnetization at the (001) termination [6]. To evaluate the robustness of our results, we carried out additional calculations using VASP with several different PBE pseudopotentials for Ru, namely Ru, Ru_pv, and Ru_sv, corresponding to the valence configurations $4d^{7}5s^{1}$, $4p^{6}4d^{7}5s^{1}$, and $4s^{2}4p^{6}4d^{7}5s^{1}$, respectively, as specified on the VASP website. All three yield the same outcome; for the unstrained (001) surface, the magnetic moments remain negligible ($10^{-6},\mu_{\mathrm{B}}$ per Ru atom). We therefore attribute the discrepancy to methodological differences, most likely related to the Brillouin-zone integration scheme, as our analysis indicates that the emergence of magnetism in RuO2 is highly sensitive to the details of $k$-space sampling.
@@
In this work, we have performed a comprehensive first-principles investigation of the magnetic properties of RuO2 thin films with (110), (100), and (001) orientations, both as freestanding slabs and in RuO2/TiO2 heterostructures. Motivated by the ongoing debate on the existence of $d$-wave altermagnetism in metallic RuO2, we systematically examined the role of epitaxial strain, surface termination, film thickness, structural relaxation, and Brillouin-zone integration schemes in determining the magnetic ground state.
@@
Our calculations demonstrate that magnetism in RuO2 thin films is intrinsically fragile and highly sensitive to both structural and methodological details. In particular, we show that the calculated magnetic moments depend strongly on the choice of Brillouin-zone integration scheme and smearing parameters. When treated with sufficiently accurate integration, bulk RuO2 at its experimental lattice parameters remains nonmagnetic, consistent with several recent experimental reports that question robust long-range antiferromagnetism in this compound. This sensitivity places RuO2 in close proximity to a magnetic instability, where small perturbations can induce local moments.
@@
Epitaxial tensile strain, mimicked by constraining RuO2 to the rutile TiO2 lattice parameters, induces finite Ru magnetic moments for all investigated surface orientations. However, in no case do we find stabilization of a fully compensated Néel-ordered AFM state and hence any altermagnetic behavior. Instead, the magnetic response is strongly layer- and site-dependent, with incomplete compensation between the two Ru sublattices (Ru1 and Ru2). In RuO2(110)/TiO2(110) heterostructures, the magnetization is dominated by surface and interface layers and exhibits pronounced oscillations with film thickness, consistent with quantum confinement and modified Ru–O hybridization at the boundaries. Even when initialized in various magnetic configurations, the system relaxes to states characterized by residual net magnetization, i.e., ferrimagnetic-like order rather than a symmetry-protected compensated AFM state.
For freestanding films, unstrained (110) and (100) slabs develop only weak surface-localized moments, while the unstrained (001) surface remains essentially nonmagnetic within numerical accuracy. Under tensile strain, all orientations develop enhanced magnetization, but the compensation between Ru1 and Ru2 moments remains incomplete. Notably, the (100) orientation exhibits the largest net magnetic moment under strain, yet still fails to realize a compensated altermagnetic phase. The (001) orientation shows the weakest magnetic response overall, even under tensile strain, making it the least favorable candidate for altermagnetic behavior among the studied surfaces.
@@
By explicitly comparing the band structures of constrained compensated AFM configurations with those of the true magnetic ground states, we show that the larger nonrelativistic spin splitting observed in thin films originates from uncompensated magnetization rather than from a symmetry-enforced altermagnetic state. Therefore, any experimentally observed spin splitting or anomalous transport in RuO2 thin films must be carefully interpreted in light of possible ferrimagnetic-like surface or interface contributions, rather than being directly attributed to ideal $d$-wave altermagnetism.
@@
Overall, our results reconcile seemingly conflicting theoretical and experimental findings by demonstrating that RuO2 resides at the brink of magnetism: small strain, reduced dimensionality, or methodological differences can induce local magnetic moments; yet none of the studied realistic thin-film geometries stabilize a fully compensated altermagnetic ground state. These findings underscore the importance of rigorous numerical convergence and detailed structural modeling when identifying candidate altermagnets, and they suggest that achieving robust altermagnetism in RuO2 may require additional symmetry engineering, chemical modification, or external tuning beyond simple epitaxial strain.
@@
In addition, our findings call for a careful re-evaluation of recent anomalous Hall measurements in RuO2 thin films that have been interpreted within a fully compensated altermagnetic framework. Since our calculations reveal an unconventional ferrimagnetic-like order with finite net magnetization in strained and heterostructured geometries, conventional Berry-curvature-driven anomalous Hall contributions are expected. These epitaxial thin films grown on TiO2 can also host strong spin-orbit couplings. We therefore propose systematic thickness- and strain-dependent transport measurements to distinguish genuine symmetry-enforced altermagnetism from the unusual uncompensated magnetism predicted in this work.

### v2

@@ III.2 RuO2(110)/TiO2(110) heterostructure
For (110) orientation, each layer contains two inequivalent Ru atoms (Ru1 and Ru2), yielding a total of six magnetic sites and ten distinct AFM configurations. We systematically examined all these AFM arrangements. Each configuration was initially imposed using magnetic constraints to stabilize the desired spin pattern; the constraints were subsequently released to allow full self-consistent relaxation. Among all tested configurations, the one initialized with antiparallel magnetic moments between Ru1 and Ru2 within each layer, i.e., a checkerboard configuration, consistently relaxes to the lowest-energy state, characterized by opposite signs of the magnetic moments on Ru1 and Ru2, with incomplete compensation. The remaining AFM initializations either converge to higher-energy solutions or evolve into the same checkerboard configuration upon relaxation. Moreover, for several slab thicknesses, calculations starting from a ferromagnetic state also converge to this checkerboard-type AFM order. These results indicate that the ground state generally favors antiparallel alignment between Ru1 and Ru2 within each layer; accordingly, this antiparallel arrangement is adopted as the initial magnetic configuration in all subsequent calculations.
@@
The only deviation from this trend occurs in the 6-L slab, where the Ru atoms in the third layer develop weak magnetic moments of the same sign on the two sublattices, despite being initialized in an AFM configuration. This behavior could be associated with small structural relaxation effects. In the third layer, the magnetic moment on Ru2 is significantly smaller than that on Ru1. The magnetic moments in the third layer are 0.33 $\mu\mathrm{B}$ on Ru1 and 0.04 $\mu\mathrm{B}$ on Ru2 for the Bridge configuration, whereas for the Cus configuration, the corresponding values are 0.30 $\mu\mathrm{B}$ and 0.007 $\mu\mathrm{B}$, respectively.
@@
Although full structural relaxations were carried out for all slab geometries, Fig. 4 shows that atomic relaxation has only a minor effect on the magnetic moments of the inner RuO2(110) layers. Moreover, when the magnetic moment of the surface layer is artificially enhanced by introducing a Hubbard $U$ correction on the Ru atoms at the surface, the induced increase remains strongly localized at the surface, while the magnetic moments in the inner layers change only negligibly. This behavior demonstrates that the magnetic moments generated at the surface due to symmetry breaking [22] do not propagate into the inner layers. Instead, the magnetic moments of the Ru atoms are primarily determined by strain-induced changes in their local atomic environment arising from interfacial tensile strain, rather than by long-range induction from the surface.
@@
Figure 3: Layer-resolved magnetic moments of Ru1 and Ru2 for film thicknesses between one and seven layers, for Bridge (top) and Cus (bottom) terminations. Filled and open symbols denote Ru1 and Ru2, respectively, for RuO2(110) thin films on a TiO2(110) substrate. Data corresponding to the same film thickness are connected by lines of the same color to guide the eye. Opposite magnetic moments of Ru1 and Ru2 atoms do not cancel each other and thus system behave like an ferrimagnet state, with varying magnetic moments at each layer, rather than a compensated AFM order in altermagnets.
@@
Figure 4: Effect of atomic relaxation and induced surface magnetization in a 5-L RuO2(110) slab on TiO2(110). Magnetic moments of Ru1 (filled symbols) and Ru2 (open symbols) are shown as a function of the Ru layer index. In the induced-magnetization scenario, we introduce a Hubbard $U$ correction on the surface Ru atoms to artificially enhance their magnetic moments. For this case, we use the geometrically relaxed structure.
To further elucidate the magnetic character, we compare the electronic structures of a hypothetically compensated AFM (altermagnetic) RuO2(110)/TiO2(110) configuration and its true ferrimagnetic-like ground state. Band structures are calculated for a five-layer RuO2(110) slab on TiO2(110) in both magnetic configurations, using the same non-relaxed geometry to exclude structural relaxation effects; see Fig. 5.
@@
To make the comparison between the two magnetic states meaningful, we adjusted the magnetic moments of the Ru atoms in the hypothetical altermagnetic case so that their total absolute magnetization matches that of the true magnetic ground state, where the magnetic moments vary from layer to layer (Fig. 3). By testing different values, we found that fixing the Ru moments to approximately $0.3~\mu_{\mathrm{B}}$ reproduces the total absolute magnetization of the unconstrained ground state. This value is therefore used in the hypothetical altermagnetic state calculations presented in the right panel of Fig. 5.
@@
As shown in Fig. 5, the band structure of the true unconstrained ground state exhibits a substantially larger nonrelativistic spin splitting than that of the hypothetical altermagnetic state. This enhancement arises because, in the true magnetic ground state, the Ru magnetic moments are not fully compensated, in contrast to the compensated AFM limit, where the net magnetization vanishes.
@@
Figure 5: Nonrelativistic band structures of a 5-L RuO2(110) slab on a TiO2(110) substrate. The left panel shows the band structure of the unconstrained magnetic ground state (the true ground state with a ferrimagnetic-like ground state), while the right panel shows the band structure obtained with Ru magnetic moments constrained to enforce a fully compensated AFM configuration (a hypothetical altermagnetic ground state). In both cases, the non-relaxed atomic structure is used. Red and blue lines denote spin-up and spin-down bands, respectively.
III.3 RuO2(100)/TiO2(100) and RuO2(001)/TiO2(001) heterostructures
@@
We next consider the remaining low-index RuO2 surfaces, namely the (100) and (001) orientations. Owing to computational limitations, only one film thickness (12-L) is examined for each case. Figure 6 shows the layer-resolved magnetic moments of the two inequivalent Ru sites, Ru1 and Ru2, before and after structural relaxation.
@@
For both orientations, atomic relaxation has a markedly stronger effect than in the (110) case. In the relaxed (001) and (100) films, magnetic moments in the outermost three layers are strongly suppressed, while inner layers near the RuO2/TiO2 interface develop larger local moments. By contrast, the unrelaxed films exhibit an nearly compensated AFM order, with Ru1 and Ru2 moments of similar magnitude and opposite sign.
@@
Despite these differences, all systems, relaxed and unrelaxed, (001), (100), and (110), exhibit a characteristic layer-by-layer oscillation in the magnitude of the local magnetic moments. This behavior can be attributed to the metallic nature of the films: electronic states are confined between the surface and the interface, forming quasi-two-dimensional states analogous to a particle-in-a-box with finite potential barriers. The resulting quantum interference gives rise to oscillatory magnetic responses across the film thickness.
@@ III.4 Intrinsic magnetism of freestanding RuO2 thin films
To further explore the role of the substrate, we also study freestanding RuO2 thin films with different surface orientations; see Fig. 1. Two classes of films are considered: (i) films constrained to the experimental lattice parameters of bulk rutile TiO2, and (ii) films constrained to the experimental lattice parameters of bulk RuO2. In both cases, only the internal atomic coordinates are relaxed, while the lattice parameters are fixed. Because the in-plane lattice constant of RuO2 exceeds that of TiO2 by approximately 0.1 Å, fixing RuO2 to the TiO2 lattice parameters introduces an effective biaxial tensile strain. This procedure, therefore, serves as a computational strategy to simulate RuO2 under tensile strain conditions.
Figure 7 shows the layer-resolved Ru magnetic moments for the (110), (100), and (001) orientations, both with and without tensile strain.
In the unstrained (110) and (100) films, finite magnetic moments are largely confined to the first two surface layers and decay rapidly toward the inner layers. The unstrained (110) result is consistent with a recent report Ref. [22]. In contrast, the (001) surface converges to a nonmagnetic solution, with no Ru atom exhibiting a discernible magnetic moment within all layers. This finding differs from a recent report of finite magnetic moments at the (001) termination [8].
To evaluate the robustness of our results in the unstrained case, we carried out additional calculations using VASP with several different PBE pseudopotentials for Ru, namely Ru, Ru_pv, and Ru_sv, corresponding to the valence configurations $4d^{7}5s^{1}$, $4p^{6}4d^{7}5s^{1}$, and $4s^{2}4p^{6}4d^{7}5s^{1}$, respectively, as specified on the VASP website. All three yield the same outcome (unshown); for the unstrained (001) surface, the magnetic moments remain negligible, $10^{-6}\mu_{\mathrm{B}}$ per Ru atom. We therefore attribute the discrepancy with Ref. [8] to methodological differences, most likely related to the Brillouin-zone integration scheme, as our analysis indicates that the emergence of magnetism in RuO2 is highly sensitive to the details of the $k$-space sampling.
@@
These results are in contrast to a recent report in Ref. [14] that found the (100) orientation to be an ideal altermagnet and the (110) orientation to be ferrimagnetic. Based on our accurate analysis, none of the investigated surfaces, (110), (100), or (001), stabilize a compensated AFM and hence an altermagnetic ground state. Again, we attribute this discrepancy to the methodology used in their ab initio calculations.
@@
In this work, we have performed a comprehensive first-principles investigation of the magnetic properties of RuO2 thin films with (110), (100), and (001) orientations, both as RuO2/TiO2 heterostructures and (un)strained freestanding slabs. Motivated by the ongoing debate on the existence of $d$-wave altermagnetism in metallic RuO2, we systematically examined the role of epitaxial strain, surface termination, film thickness, structural relaxation, and Brillouin-zone integration schemes in determining the magnetic ground state.
@@
Our calculations demonstrate that the true magnetic ground state in RuO2 thin films is intrinsically fragile and highly sensitive to both structural and methodological details. In particular, we show that the calculated magnetic moments depend strongly on the choice of the Brillouin-zone integration scheme and smearing parameters. When treated with sufficiently accurate integration, bulk RuO2 at its experimental lattice parameters remains nonmagnetic, consistent with several recent experimental reports that question robust long-range antiferromagnetism in this compound. This sensitivity places RuO2 in close proximity to a magnetic instability, where small perturbations can induce local moments.
@@
In RuO2(110)/TiO2(110) heterostructures, the magnetization is dominated by surface and interface layers and exhibits pronounced oscillations with film thickness, consistent with quantum confinement and modified Ru–O hybridization at the boundaries. Even when initialized in various magnetic configurations, the system relaxes to states characterized by residual net magnetization, i.e., an unusual ferrimagnetic-like order rather than a compensated AFM order with altermagnetic behavior.
For freestanding and unstrained (110) and (100) slabs, only weak surface-localized moments develop, while the unstrained (001) surface remains essentially nonmagnetic within numerical accuracy.
Tensile biaxial strain, mimicked by constraining freestanding RuO2 thin film to the rutile TiO2 lattice parameters, induces finite Ru magnetic moments for all investigated surface orientations. However, in no case do we find stabilization of a compensated Néel-ordered AFM order and hence any altermagnetic behavior. The (100) orientation exhibits the largest net magnetic moment under strain, while the (001) orientation shows the weakest magnetic response overall.
@@
Through a direct comparison between a hypothetical compensated AFM configuration, i.e., altermagnet state, and the true magnetic ground state, i.e., ferrimagnetic-like state, we show that the pronounced spin splitting in thin films stems from the uncompensated magnetization, not from symmetry-enforced altermagnetism. Therefore, experimentally observed spin splitting or anomalous transport in RuO2 thin films should not be uncritically attributed to $d$-wave altermagnetism, but instead evaluated in view of possible emerging unconventional ferrimagnetic-like states in the system.
@@
Overall, our results reconcile seemingly conflicting various theoretical and experimental findings by demonstrating that RuO2 resides at the brink of magnetism: methodological differences, small strain, or reduced dimensionality, can induce small local magnetic moments; yet none of the studied realistic thin-film geometries stabilize a compensated altermagnetic ground state. These findings underscore the importance of rigorous numerical convergence and detailed structural modeling when identifying candidate altermagnets, and they suggest that achieving robust altermagnetism in RuO2 may require additional symmetry engineering, chemical modification, or external tuning beyond simple epitaxial strain.
@@
In addition, our findings call for a careful re-evaluation of recent anomalous Hall measurements in RuO2 thin films that have been interpreted within a fully compensated altermagnetic framework. Since our calculations reveal an unconventional ferrimagnetic-like state with finite net magnetization in epitaxial RuO2 thin films, implying conventional Berry-curvature-driven anomalous Hall effects. Given the strong spin–orbit coupling in RuO2/TiO2 heterostructures, we propose systematic thickness- and strain-dependent transport measurements to distinguish genuine symmetry-enforced altermagnetism from the uncompensated magnetism predicted here.
V Method
We performed nonrelativistic, ab initio density functional theory (DFT) calculations using the Quantum ESPRESSO (QE) package [16, 15, 17]. The electron–ion interactions were described using the Standard Solid-State Pseudopotentials (SSSP) library, PBEsol Efficiency version 1.3.0 [45]. For the exchange–correlation functional, we employed the PBEsol approximation [43], a revised form of the Perdew–Burke–Ernzerhof (PBE) generalized gradient approximation (GGA) that provides improved accuracy for the structural properties of solids and surfaces. A vacuum spacing of approximately 15 Å was introduced in all thin-film calculations to avoid spurious interactions between periodic images. Brillouin-zone integrations were performed using a uniform $k$-point grid corresponding to a reciprocal-space resolution of 0.10 Å-1. The thin-film structures were constructed using the CIF2Cell code [7], and the initial QE input files were generated with the QE input generator [45]. The plane-wave kinetic energy cutoff was set to 40 Ry for the wave functions and 400 Ry for the charge density. The magnetic moment associated with each Ru atom was obtained using Bader charge analysis as implemented in the Critic2 code [40, 41].
Data availability
The data that support the findings of this study are available upon reasonable request.
Author Contributions
M.A. and A.Q. jointly defined the project. M.A. conceived and supervised the research, performed the majority of the DFT calculations, and co-wrote the manuscript. N.R. carried out structural visualizations, contributed to selected DFT calculations, and provided intellectual input. I.M. performed additional simulations. A.R.O. provided intellectual input and financial support. M.A. and A.Q. jointly discussed the results and finalized the manuscript. All authors reviewed and edited the manuscript.
Competing interests
The authors declare no competing interests